8) Write the first four terms of a Maclaurin series for 𝑓(đ‘„)=đ‘’đ‘„?

1
Expert's answer
2022-04-13T12:13:59-0400

The Taylor series of a function f (x) that is infinitely differentiable at a number "a"  is the power series

"f(x)=f(a)+\frac{f'(a)}{1!}(x-a)+\frac{f''(a)}{2!}(x-a)^2+""\frac{f'''(a)}{3!}(x-a)^2+..." .

When "a=0", the series is also called a Maclaurin series.

"f(x)=e^x" , "f(0)=1" ;

"f'(x)=e^x" , "f'(0)=1" ;

"f''(x)=e^x" , "f''(0)=1" ;

"f'''(x)=e^x" , "f'''(0)=1" .

"e^x=1+\frac{1}{1!}(x-0)+\frac{1}{2!}(x-0)^2+\frac{1}{3!}(x-0)^3+...=""1+x+\frac12x^2+\frac16x^3+..."

Answer: "e^x=1+x+\frac12x^2+\frac16x^3+..." .